I do not understand inverse and direct variation.
How would i set this up?
suppose z varies directly as x and inversly as the square of y. If x=5 and y=2, then z=40. find z when x=6 and y=4

Answer by stanbon(75887)

suppose z varies directly as x and inversly as the square of y.
z = k*x/y^2
---------------
Solve for "k" using "If x=5 and y=2, then z=40".
40 = k*5/(2^2)
160 = 5k
k = 160/5
-----------------
Equation:
z = (160/5)x/y^2
------------------------
find z when x=6 and y=4
z = (160/5)*6/16
---
z = 10*6/5
z = 12
===============
Cheers,
Stan H.
================
-------